# Schemex: Interactive Structural Abstraction from Examples with Contrastive Refinement

### v1

@@ Abstract.
Figure 1. Schemex is an interactive, AI-powered visual workflow that helps users induce schemas from examples. It comprises three main stages: clustering, abstraction, and contrastive refinement.
@@ 1. Introduction
All creative or communicative works are underpinned by structural frameworks. When people undertake tasks such as writing a research abstract, designing a slide deck, or making a product demo, they are rarely starting from a blank slate. Instead, they draw on an internalized understanding of how such artifacts are typically organized—what components are expected, how they are sequenced, and what makes the result coherent and effective. This structural knowledge, while often tacit rather than explicit, plays a critical role in both interpreting existing examples and producing new ones. By examining multiple instances of a genre or task, people begin to identify recurring patterns—mental templates that organize content and guide its construction. The clearer and more explicit this structure becomes, the better it supports purposeful, flexible creation rather than mere surface-level imitation. In cognitive science, such internalized structural representations are known as schemas: abstract models that capture the essential components of a concept and the relationships among them.
Schema induction is the process of discovering the abstract structure from unstructured examples. Experts often create schemas implicitly from years of analyzing and creating examples. However, schemas can also be made explicit—these explicit schemas (also known as design patterns) are extremely helpful to novices in structuring and guiding their creative process. The cognitive process of schema induction broadly resembles the iterative process of sensemaking, incorporating both inductive and deductive approaches. One must iteratively verify that the abstractions and structures induced from the examples actually fit those examples, and when discrepancies arise, deduce how to reconcile them. For example, from looking at examples of HCI papers, one might learn that the first sentence is typically broad motivation, meant to show why the problem or areas the paper addresses are important. By looking at multiple examples, the recurring patterns—and variations within them, can be deduced.
@@
To address these challenges, we frame schema induction as an interactive process grounded in real-world examples and guided by AI-assisted pattern discovery. Drawing on recent advances in AI reasoning—such as DeepSeek R1’s “slow thinking” capabilities (Guo et al., 2025)—we explore how AI can scaffold the core cognitive steps of clustering, abstraction, and contrastive refinement.
@@
  • Clustering: Starting from a set of examples, the AI identifies latent groupings based on structural similarities and divides examples into groups.
@@
  • Abstraction: Within each cluster, it infers underlying dimensions and derives both overall and dimension-specific attributes.
  • Contrastive Refinement: The current schema is used to generate outputs, which are then compared to real (“gold”) examples and iterate the schema. To guard against overfitting, unseen validation examples are also included.
Throughout the process, humans play a central role as reflective evaluators—they make critical judgments about coherence and utility, direct the exploration process, and retain decision agency over when to advance or refine the schema. This approach grounds schema induction in concrete examples, promoting transparency, adaptability, and effective human-AI collaboration.
@@
To operationalize this approach, we developed Schemex, an interactive visual workflow for schema induction. Schemex scaffolds interactive structural abstraction and constrastive refinement through AI-assisted reasoning and visualizations that keep users grounded in real examples. Users initiate the process by specifying a goal (e.g., “Write HCI paper abstracts”) and providing a set of examples (e.g., CHI best paper abstracts). These examples can be multimodal; Schemex employs models such as GPT-4V and Whisper to convert images or videos into structured textual descriptions. The system clusters examples based on structural similarity and displays them as interactive visual nodes. Users can explore these clusters to inspect shared features and example-level mappings, validating the coherence of the clustering. They can then select a cluster of interest (e.g., empirical studies) and prompt the system to extract key structural dimensions (e.g., Motivation, Method, Findings). A visualized matrix allows users to examine how strongly each example reflects each dimension, supported by in-line explanations and citations. Users guide the next step by requesting inferences of dimension-level attributes (e.g., “Challenge Statement” under Motivation) and overall attributes (e.g., tone, length, or style). They evaluate the coherence of these attributes across examples and explore implementation details through contextual snippets. Users then apply the derived schema: given an input (e.g., a paper title), the system generates content guided by the schema. Schemex enables users to compare generated outputs and original examples via a color-coded dimensional analysis interface. Differences are highlighted, and the system suggests schema refinements that can be incorporated in the iteration. Validation cases not used in earlier steps are also provided, allowing further testing of schema robustness. This iterative cycle—observe, apply, compare, refine—continues until users arrive at a coherent, actionable schema tailored to their goal. In our user study, participants reported significantly greater insight and confidence in the schemas developed with Schemex compared to those created with the baseline condition, OpenAI o1-pro. Moreover, both schemas and outputs generated with Schemex were rated significantly higher in quality and depth by expert evaluators compared to those produced using the baseline condition.
@@
  • A structural abstraction technique that leverages clustering and pattern analysis to infer dimensions and attributes from example sets, including multimodal artifacts;
  • A contrastive refinement method that supports structured generation tasks by comparing AI-generated and real examples to guide schema evolution;
  • An interactive visual workflow that links inferred schema elements to source examples, enabling interpretable and revisable schema induction;
  • An empirical evaluation showing that schemas and outputs produced with Schemex are rated significantly higher in quality and example alignment than those from a strong AI baseline (o1-pro).
@@ 2.1. Schema Induction and Cognitive Foundations
Humans can infer schemas from concrete experiences. Cognitive science theories of analogy and structural alignment suggest that comparing multiple examples enables people to abstract their common relational structure (Gick and Holyoak, 1983; Gentner and Markman, 1997). In classic analogical problem-solving studies, Gick and Holyoak (Gick and Holyoak, 1983) demonstrated that examining two analogs can lead to the induction of a schema—a generalized solution framework that preserves shared relations while ignoring surface differences. Such analogical schemas support the transfer of knowledge to novel problems. These cognitive theories provide a foundation for the idea that people can induce schemas or conceptual models by observing recurring patterns across multiple instances. An effective schema abstracts away surface details to highlight underlying structure. For example, prior work on schema induction (Yu et al., 2014a) mentions that, to capture the essence of an invention, it is more useful to focus on its purpose (e.g., detaching components) and mechanism (e.g., using comb-like features), rather than vague notions (e.g., “making things easy”) or overly specific traits (e.g., color). However, striking the right balance between specificity and generalizability remains a challenge in schema induction.
@@ 2.2. Approaches of Interactive Structural Abstraction and Contrastive Refinement
A complementary technique for refining abstractions is contrastive refinement—clarifying boundaries by comparing positive, negative, or near-miss examples. MOCHA (Gebreegziabher et al., 2024) illustrates this by generating counterexamples that expose the limits of a user’s current concept, prompting structural correction and alignment. Flash Fill (Gulwani, 2011) supports refinement by allowing users to correct inferred transformation rules using contrasting input-output examples. These systems show how surfacing edge cases and exceptions can accelerate convergence toward accurate and generalizable models.
Schemex builds on these traditions of mixed-initiative abstraction and contrastive refinement but targets multimodal design examples and schema-level generalization. Schemex treats structure learning as a collaborative, iterative process. It specifically supports users in moving from scattered examples to coherent, named schemas through clustering, structural abstraction, and contrastive refinement. This workflow makes schema induction more interpretable and actionable for downstream design and authoring tasks.
@@ 2.3. HCI Systems for Structure Discovery and Sensemaking
Schemex builds on a growing body of HCI research that supports structure discovery and sensemaking, particularly in design and multimodal content domains. In design, many systems aim to surface reusable patterns by mining common structures at scale. Prior work has emphasized documenting reusable solutions in structured schema formats (e.g., problem, context, solution) (Kruschitz and Hitz, 2010). For example, Webzeitgeist mined common UI structures from thousands of websites (Kumar et al., 2013), and other systems have identified interface design patterns across mobile apps through large-scale clustering (Alharbi and Yeh, 2015). Recent systems such as DesignWeaver (Tao et al., 2025) and Luminate (Suh et al., 2024) help users explore the design space of AI-generated variations by revealing latent design dimensions and attributes.
@@
Schemex shares the goal of supporting structure discovery through mixed-initiative workflows, targeting schema induction from multimodal content creation examples. Unlike prior systems that focus on AI-generated artifacts (e.g.,  (Suh et al., 2024; Tao et al., 2025)), Schemex emphasizes human sensemaking from real artifacts. It supports users in articulating and evolving actionable schemas. In this way, Schemex complements and extends prior systems by offering a workflow for interpretable, interactive, and grounded schema authoring.
@@ 2.4. Schemas Used in HCI Systems
Because of the generalized scaffolding schemas provide—particularly for novices—schemas are widely used in creativity support tools. Books of design patterns have influenced generations of architects (Alexander et al., 1977), software engineers (Gamma et al., 1995), and web designers (Duyne et al., 2002). In computer graphics, design patterns can help solve challenges such as laying out furniture in a room (Merrell et al., 2011), generating usable maps (Agrawala and Stolte, 2001), illustrating furniture assembly instructions (Agrawala et al., 2011), or sequencing cuts in film (Leake et al., 2017). In HCI, design patterns have been used to help novice filmmakers structure their videos (Kim et al., 2015), support human-robot interaction programming (Sauppé and Mutlu, 2014; Kahn et al., 2008), and inspire novel product ideas (Yu et al., 2014a, b). AI has also been instrumental in helping users apply abstract schemas to creative tasks such as writing stories (Mirowski et al., 2023), creating visual blends (Chilton et al., 2019, 2021; Cunha et al., 2017, 2018; Wang et al., 2023), producing trailers and teasers (Wang et al., 2024b), crafting social media videos (Wang et al., 2024a; Menon et al., 2024), and generating inclusive, persuasive emergency messages (Jit et al., 2024). In nearly all of these systems, authors have had to manually analyze examples and construct schemas by hand—a labor-intensive process. Accelerating schema induction could vastly expand access to structured creativity, enabling more people to design, communicate, and express ideas effectively.

### v2

@@ Abstract.
Figure 1. Schemex helps novices induce schemas from real-world examples (e.g., CHI paper abstracts). (1) Input: users provide an example set. (2) Clustering: the system groups similar examples and displays a feature×example matrix so users can inspect and adjust clusters. (3) Abstraction: for a selected cluster, Schemex proposes a candidate schema: key dimensions and attributes (plus overall attributes) and populates an example×dimension/attribute matrix; users can add, rename, or remove elements. (4) Contrastive refinement loop: the current schema is applied to generate outputs, which are compared against gold examples, color-coded by dimensions. Common gaps and improvement suggestions are highlighted. Users review these suggestions, and the system incorporates approved changes to refine the schema. The loop continues until arriving at a coherent, actionable schema that is tailored to the user’s goal.
@@ 1. Introduction
All creative or communicative works are underpinned by structural frameworks (Campbell, 2008; Gamma et al., 1995; Fox and Weissman, 2013). When people undertake tasks such as writing a research abstract, designing a slide deck, or making a product demo, they are rarely starting from a blank slate. Instead, they draw on an internalized understanding of how such artifacts are typically organized—what components are expected, how they are sequenced, and what makes the result coherent and effective. This structural knowledge, while often tacit rather than explicit, plays a critical role in both interpreting existing examples and producing new ones. By examining multiple instances of a genre or task, people begin to identify recurring patterns—mental templates that organize content and guide its construction. The clearer and more explicit this structure becomes, the better it supports purposeful, flexible creation rather than mere surface-level imitation. In cognitive science, such internalized structural representations are known as schemas (Gick and Holyoak, 1983): abstract models that capture the essential components of a concept and the relationships among them.
Schema induction is the process of discovering abstract structure from unstructured examples. Experts often develop schemas implicitly through years of analyzing and creating examples. However, schemas can also be made explicit—these explicit schemas (also known as design patterns (Gamma et al., 1995)) provide reusable frameworks that help people organize complex tasks, scaffold ideation, and transfer knowledge across domains, making them especially valuable for novices. The cognitive process of schema induction resembles the iterative cycle of sensemaking (Pirolli and Card, 2005): people induce candidate structures from recurring patterns across examples, and then deduce whether these structures hold by testing them against further instances. This back-and-forth ensures that abstractions remain grounded in reality while still capturing generalizable insights. For example, from analyzing multiple HCI abstracts, one might induce that opening sentences establish broad motivation. By deducing how this pattern applies across different papers—say, a system paper vs. a field study—one can refine the schema to capture both shared conventions and genre-specific variations.
@@
We frame schema induction as an iterative process of hypothesizing a structure, applying it to examples, and refining it by identifying where the schema fails. Our approach follows a three-stage workflow that mirrors the cognitive steps of schema formation while directly targeting the challenges:
@@
  • Clustering: When examples differ structurally, forcing them into a single schema produces weak generalizations. To avoid this, AI identifies latent groupings based on structural similarities and divides examples into groups.
@@
  • Abstraction: Within each cluster, it infers underlying dimensions and derives both overall and dimension-specific attributes. This synthesizes multiple examples into a schema that captures generalizable patterns while easing the cognitive demand of making sense of many examples.
  • Contrastive Refinement: Schemas cannot be judged abstractly; they must be tested in use. Contrastive refinement applies candidate schemas, compares generations to real examples, and iterates until schemas are precise and actionable. To guard against overfitting, unseen validation examples are also included.
Humans remain central throughout as reflective evaluators: they validate clusters, interpret abstractions, and decide how to refine schemas. AI serves as a scaffold that surfaces hidden patterns and accelerates iteration, while preserving human agency in shaping the final schema.
@@
To operationalize this approach, we developed Schemex, an interactive visual workflow for schema induction. Schemex scaffolds interactive structural abstraction and contrastive refinement through AI-assisted reasoning and visualizations that keep users grounded in real examples. Users initiate the process by specifying a goal (e.g., “Write HCI paper abstracts”) and providing a set of examples (e.g., CHI best paper abstracts). These examples can be multimodal; Schemex employs models such as GPT-4V and Whisper to convert images or videos into structured textual descriptions. The system clusters examples based on structural similarity and displays them as interactive visual nodes. Users can explore these clusters to inspect shared features and example-level mappings, validating the coherence of the clustering. They can then select a cluster of interest (e.g., empirical studies) and prompt the system to extract key structural dimensions (e.g., Motivation, Method, Findings). A visualized matrix allows users to examine how strongly each example reflects each dimension, supported by in-line explanations and citations. Users guide the next step by requesting inferences of dimension-level attributes (e.g., “Qualitative insights” under Findings) and overall attributes (e.g., tone, length, or style). They evaluate the coherence of these attributes across examples and explore implementation details through contextual snippets. Users then apply the derived schema: given an input (e.g., a paper title), the system generates content guided by the schema. Schemex enables users to compare generated outputs and original examples via a color-coded dimensional analysis interface. Differences are highlighted, and the system suggests schema refinements that can be incorporated in the iteration. Validation cases not used in earlier steps are also provided, allowing further testing of schema robustness. This iterative cycle—observe, apply, compare, refine—continues until arriving at a coherent, actionable schema tailored to the user’s goal.
@@
  • A new technique of contrastive refinement that supports structured generation tasks by applying schemas to diverse cases, comparing them with real examples, and guiding iterative evolution;
  • An interactive visual workflow that integrates AI reasoning with example-based evidence, enabling users to explore, inspect, and revise schemas through linked visualizations;
  • A technical evaluation demonstrating that Schemex produces significantly higher-quality schemas and outputs than a strong AI baseline (o1-pro);
  • A user study showing that Schemex’s interactive process fosters deeper learning in schema induction, with participants reporting a 147% increase in structural insight and a 148% increase in novel reflection.
@@ 2.1. Schema Induction and Cognitive Foundations
Humans can infer schemas from concrete experiences. Cognitive science theories of analogy and structural alignment suggest that comparing multiple examples enables people to abstract their common relational structure (Gick and Holyoak, 1983; Gentner and Markman, 1997). In classic analogical problem-solving studies, Gick and Holyoak (Gick and Holyoak, 1983) demonstrated that examining two analogs can lead to the induction of a schema—a generalized solution framework that preserves shared relations while ignoring surface differences. Such analogical schemas support the transfer of knowledge to novel problems. An effective schema abstracts away surface details to highlight underlying structure. For example, prior work on schema induction (Yu et al., 2014a) mentions that, to capture the essence of an invention, it is more useful to focus on its purpose (e.g., detaching components) and mechanism (e.g., using comb-like features), rather than vague notions (e.g., “making things easy”) or overly specific traits (e.g., color of the invention). However, striking the right balance between specificity and generalizability remains a challenge in schema induction.
@@ 2.2. Approaches of Interactive Structural Abstraction and Contrastive Refinement
A complementary strategy is to use contrast to sharpen abstractions—typically as contrastive learning or evaluation—by comparing positive, negative, or near-miss examples. Mocha (Gebreegziabher et al., 2025), for instance, generates counterexamples that are syntactically and semantically similar but predicted with different labels. By asking users to annotate these counterexamples, Mocha improves model learning while helping users reflect on and refine their concept definitions. Similarly, Flash Fill (Gulwani, 2011) lets users refine inferred transformation rules via contrasting input–output pairs. These systems show how surfacing edge cases and exceptions can accelerate convergence toward accurate and generalizable models.
Schemex builds on these abstraction and contrastive techniques but extends them to schema-level generalization over real design examples. Unlike systems such as Mocha that use counterexamples for concept labeling tasks, Schemex introduces the contrastive refinement technique that applies candidate schemas, compares outputs against real examples along named dimensions, and revises schemas based on diagnosed mismatches. This process turns contrasts into actionable edits of schema structures, enabling users to move from scattered examples to coherent, named schemas that are interpretable and actionable for downstream design and authoring tasks.
@@ 2.3. HCI Systems for Structure Discovery and Sensemaking
Schemex builds on a growing body of HCI research that supports structure discovery and sensemaking, particularly in design and content creation domains. In design, many systems aim to surface reusable patterns by mining common structures at scale. Prior work has emphasized documenting reusable solutions in structured schema formats (e.g., problem, context, solution) (Kruschitz and Hitz, 2010). For example, Webzeitgeist mined common UI structures from thousands of websites (Kumar et al., 2013), and other systems have identified interface design patterns across mobile apps through large-scale clustering (Alharbi and Yeh, 2015). Recent systems such as DesignWeaver (Tao et al., 2025) and Luminate (Suh et al., 2024) help users explore the design space of AI-generated variations by revealing latent design dimensions and attributes.
@@
Schemex shares the goal of supporting structure discovery through mixed-initiative workflows, but differs in scope and emphasis. Unlike systems such as LLooM (Lam et al., 2024) that focus on clustering or induction, Schemex supports the more demanding task of authoring complete schemas. And unlike tools like DesignWeaver (Tao et al., 2025) and Luminate (Suh et al., 2024) that analyze AI outputs from known prompts, Schemex addresses the harder problem of discovering new, verified schema knowledge from messy, real-world examples that extend beyond what LLMs already encode. In this way, Schemex complements and extends prior systems by offering a workflow for interactive and grounded schema authoring.
@@ 2.4. Schemas Used in HCI Systems
Because of the generalized scaffolding schemas provide—particularly for novices—schemas are widely used in creativity support tools. Books of design patterns have influenced generations of architects (Alexander et al., 1977), software engineers (Gamma et al., 1995), and web designers (Duyne et al., 2002). In computer graphics, design patterns can help solve challenges such as laying out furniture in a room (Merrell et al., 2011), generating usable maps (Agrawala and Stolte, 2001), illustrating furniture assembly instructions (Agrawala et al., 2011), or sequencing cuts in film (Leake et al., 2017). In HCI, design patterns have been used to help novice filmmakers structure their videos (Kim et al., 2015), support human-robot interaction programming (Sauppé and Mutlu, 2014; Kahn et al., 2008), and inspire novel product ideas (Yu et al., 2014a, b). AI has also been instrumental in helping users apply abstract schemas to creative tasks such as writing stories (Mirowski et al., 2023), creating visual blends (Chilton et al., 2019, 2021; Cunha et al., 2018; Wang et al., 2023), producing trailers and teasers (Wang et al., 2024b), crafting social media videos (Wang et al., 2024a; Menon et al., 2024), and generating inclusive, persuasive emergency messages (Jit et al., 2024). In nearly all of these systems, authors have had to manually analyze examples and construct schemas by hand—a labor-intensive process. Accelerating schema induction could vastly expand access to structured creativity, enabling more people to design, communicate, and express ideas effectively.
@@ 3. Schemex System
Inducing schemas from examples is valuable but challenging, especially for novices. Surface variations can obscure underlying structural similarities, examples within a single set may demand different treatments, and the adequacy of a schema is often revealed only when tested across cases. Effective schemas must balance specificity—providing enough guidance for action—with generality—supporting transfer to new contexts.
To address these challenges, Schemex (see Figure 2) supports an iterative, example-grounded workflow that keeps abstractions accountable to source evidence and easy to refine. Its design draws from three foundations: (a) sensemaking theories emphasizing the construction and revision of representational artifacts (Pirolli and Card, 2005), (b) research privileging relational structure over surface form (Gentner, 1983), and (c) Human-AI interaction guidelines that recommend exposing system state and rationale while giving users meaningful control (Settles, 2009).
This section introduces the design goals of Schemex, presents a motivating user walkthrough, describes its key features, and concludes with implementation details.
3.1. Design Goals
Schemex was developed to support novices learning an unfamiliar genre or form who seek to infer underlying schemas from real-world examples. The system is designed to foster reflective schema induction, rather than generic content generation. Its design is guided by four goals:
  1. DG1
    Ensure schemas remain grounded in evidence. Schema induction must remain accountable to the source examples. Schemex ensures that both AI-generated proposals and human judgments are anchored in evidence through inline citations, matrices, and rationales. This groundedness supports trust, transparency, and alignment with established approaches (Chilton et al., 2013).
  2. DG2
    Ensure examples share structural similarities before developing schemas. Generalizing across structurally different examples can produce weak or overly broad schemas. Clustering provides a safeguard by grouping examples that share latent similarities. This prevents overgeneralization and establishes a basis for abstraction, consistent with sensemaking accounts that stress organizing artifacts to reveal structure (Pirolli and Card, 2005; Gentner, 1983).
  3. DG3
    Evaluate schema quality by applying it to examples. Schema quality cannot be judged abstractly—it becomes visible only when applied across diverse cases. Schemex operationalizes this principle by applying schemas, inspecting contrasts, and revising iteratively. This apply-and-test cycle reflects research on contrastive learning and near-boundary reasoning (Settles, 2009).
  4. DG4
    Allow human interactive editing of schemas. While AI can propose clusters and structures, it lacks contextual knowledge of the user’s goals and domain. Effective schema induction therefore requires human oversight at critical junctures: to validate clusters, accept or reject structures, and steer revisions. Schemex provides lightweight, localized controls to preserve agency while keeping iteration fast and cognitively manageable (Amershi et al., 2019; Sweller, 1988).
3.2. Motivating User Walkthrough
Consider Sam, a first-year graduate student learning to write HCI abstracts. Sam begins by entering the creation goal—“Write an HCI paper abstract”—and uploads 20 CHI’24 Best Paper abstracts along with their titles. The system stores each abstract–title pair and automatically sets aside a small validation subset for later testing.
3.2.1. Stage 1: Clustering.
@@ 3.2. Motivating User Walkthrough
Sam clicks Run Clustering (see Figure 3). Schemex processes all examples at once and proposes three clusters: Empirical Studies, Theoretical Contributions, and System Design and Evaluation. Each cluster is represented as a node on the interface. Curious about theory papers, Sam clicks on the Theoretical Contributions node. A side panel opens, showing common features shared by the cluster: 1) Discussion of frameworks or concepts that advance understanding; 2) Emphasis on implications for theory or policy; 3) Engagement with existing literature to reframe or extend prior theories.
Below this summary, a feature-example relationship matrix lists each abstract in the cluster and indicates (yes/partial/no) whether it exhibits each feature, with citations to the supporting text spans. For example, the snippet “…generates implications for design and policy changes” shows that the paper abstract is about emphasis on implications for theory or policy. Sam skims the side panel, learns the common features of each cluster, and confirms that the clustering is coherent, deciding to proceed with this subset. This ensures schema induction is grounded in evidence (DG1) and prevents overgeneralization across mismatched examples (DG2).
3.2.2. Stage 2: Abstraction.